OpenStudy (anonymous):

a body is thrown from a tower of height 45m another body is thrown up at the same instant with speed u from a height of 30m if they reach the ground simultaneously , u is(g=10m/s^2)

OpenStudy (anonymous):

the ans=5m/s

OpenStudy (ajprincess):

For the first body 45=1/2*10*t^2 45=5t^2 t^2=9 t=3 For the second body -30=3u-1/2*10*9 -30=3u-45 45-30=3u 15=3u u=5m/s